Position of one particle given by x=2t^2-t^3 so find the distance travelled by the particle in first 3 seconds.

Grade:12th pass

1 Answers

Arya
5 years ago
TO GET THE POSITION OF THE PARTICLE AT T=3 PUT T=3 IN THE EQUATION 
2(3)2-(3)3
=2*9-27
=18-27=-9M
ANS.-9M